OpenStudy (anonymous):

Part 1: What are the possible number of positive, negative, and complex zeros of f(x) = –2x3 – 5x2 + 6x + 4? Part 2: Use complete sentences to explain the method used to solve this equation.

OpenStudy (anonymous):

you can Use Descartes' Rules of Signs

OpenStudy (anonymous):

positive is 1

OpenStudy (anonymous):

not sure how though/.. can u explain?

OpenStudy (anonymous):

to find the possible positive number look at the equation where the term change its sign in the equation -5x^2 has negative sign the next term to it is 6x which is positive , therefore in the equation there is one only that change its sign,, from negative to positive

OpenStudy (anonymous):

@sar12 that is deacartes rule

OpenStudy (anonymous):

right i got that so what else?

OpenStudy (anonymous):

well so -2x3 goes to -5x2 which is not one?

OpenStudy (anonymous):

so only one P, because of -5x^2 to +6x?

OpenStudy (anonymous):

so u just change the opposite signs of everything in the equation for the negative?

OpenStudy (anonymous):

yes good, for negative just change f(x) to f(-x) and find how many change its terms, like this [f(-x) = -2(-x)^3 -5(-x)^2 + 6(-x) + 4

OpenStudy (anonymous):

the x to -x

OpenStudy (anonymous):

so now we have 1 Positive, 1 negative? Why dont we have 1 or 0? what does that have to do with anything anyway?

OpenStudy (anonymous):

@sar12 so what did you get in negative???

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

how i thought we are counting the sign changes?

OpenStudy (anonymous):

yup im sorry its 1 xD

OpenStudy (anonymous):

its ok, so why sometimes its like also 0 like 2 and 0?

OpenStudy (anonymous):

http://www.purplemath.com/modules/drofsign.htm to understand the whole concept of Descartes' Rules of Signs I think this will help you a lot

OpenStudy (anonymous):

ok thanks wait so what about the complex?

OpenStudy (anonymous):

@traile what about the complex?

OpenStudy (anonymous):

um so do u know? can u help me?

OpenStudy (anonymous):

wahts a complex zero? I need help with that ....

OpenStudy (anonymous):

get the roots of the equation the answer is 2

OpenStudy (anonymous):

2 complex zeros, equate the equation to zero then get the roots

OpenStudy (anonymous):

get the roots, what is the roots? how to do u evaluate?

OpenStudy (anonymous):

is it two or zero?

OpenStudy (anonymous):

like Complex zeros: 2, 0 ?

OpenStudy (anonymous):

Im not that sure but I think the answer is 2 or 0, Complex 2 or 0

OpenStudy (anonymous):

but dont u have to subtract by 2 if its like 2 or more?,

OpenStudy (anonymous):

For example if there were two positive, then it would be 2 or 0...

OpenStudy (anonymous):

so wait its the thrid degree so that is where u find the complex roots? how do u find the complex zeros?

OpenStudy (campbell_st):

A cubic, will have 3 roots, they may be repeated, complex, irrational or rational. there is a root between x = -1 and x = 0 using the remainder theorem P(-1) = 2 - 5 -6 + 4 = -5 P(0) = 4 try P(-1/2) = 1/4 - 5/4 - 6/2 + 4 P(1/2) = 0 so x = -1/2 is a root or (2x +1) is a factor use polynomial division to find the quadratic solution  so\[ P(x) = (2x+1)(-x^2 -2x + 4)\] you will need the general quadratic formula for the roots of the quadratic